OpenStudy (anonymous):

Write in radical notation. evaluate the radical expression. Write the expression in radical notation

OpenStudy (anonymous):

Write the expression 512 ^1/3 in radical notation

OpenStudy (anonymous):

\[x ^{1/2}=\sqrt{x}\]

OpenStudy (anonymous):

was that for 512^1/3

OpenStudy (anonymous):

yea its the same concept. \[x ^{1/3}=\sqrt[3]{x}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem, do you understand?

OpenStudy (anonymous):

I wrote my problem incorrectly, just got back from dinner. fresh eyes. lol it says write the expression 512^-1/3 in radical notation.

OpenStudy (anonymous):

do you know what a negative exponent means?